# Illusion or Algorithm? Investigating Memorization, Emergence, and Symbolic Processing in In-Context Learning

### v2

@@ 3 Experimental Setup: Tasks, Data and Models
Previous literature primarily defines ICL as the ability to learn an input-label map based on the explicit samples provided in-context, with only exception being Elhage et al. (2021), who investigated literal sequence copying as a precursor to ICL. In this work, we argue that any next-token prediction that requires the model to recover a token distribution from the context and utilize it (with or without another token distribution memorized from the pre-training) qualifies as ICL—a notion of ICL very similar to Xie et al. (2022). This definition allows us to define pattern-matching tasks similar to Olsson et al. (2022), standard ICL examples of input-label pairs, as well as counterfactual reasoning with in-context facts under the broad umbrella of ICL.
@@
To the best of our knowledge, all prior works evaluating ICL (Conneau & Kiela, 2018; Hernandez et al., 2024; Todd et al., 2023) have used tasks that require some form of factual, common-sense, or linguistic knowledge and therefore may be subject to data contamination. Since these web-scale training datasets often comprise billions or even trillions of tokens, we cannot completely rule out the possibility—even when the pre-training dataset is fully accessible (e.g., Biderman et al., 2023; Groeneveld et al., 2024). Additionally, with real-world, natural-language tasks, it is practically impossible to control task configurations (e.g., feature sparsity, length dependence, token frequency statistics, reliance on parametrized knowledge, etc.), and as a result, one cannot reliably correlate model scale (training time and parameter size) with specific task characteristics.
Proposed task suit.
To address the limitations of using existing benchmark data as explained earlier, in Table 1, we propose a novel suite of tasks constructed from various patterns using randomly sampled tokens, including literal sequence copying (lsc), literal sequence copying with random gaps (lscg), word content (wc) and word index (wi). These tasks satisfy the notion of ICL as recovering and inferring from token distributions in-context. Since these tasks appear as sequences of random tokens, it is functionally impossible for them to exist in the pre-training dataset. Therefore, if an LM can perform these tasks, it cannot be due to memorization—some form of pattern generalization must be at play. Moreover, we introduce two tasks focused on associating and overriding parameterized information using the ICL pattern identification capability: we use token translation [^2] For the tt task, we sample 200 common words across four languages: English, German, Spanish, and Italian. The sampling process is detailed in Appendix A. (tt) to investigate how patterns are combined with parameterized information, and counterfactuals (cf) to observe how this parameterized information can be overridden by counterfactual prompt settings. Since these tasks require specific parameterized information, we cannot guarantee that it does not appear in the pre-training dataset; however, this is acceptable, as we already include pattern-based tasks that do not rely on such information and serve as a cleaner test of ICL capabilities. We assume that they are “harder” than the pattern-based tasks, as they involve an additional layer of combining or overriding information.
@@
Table 2: Selected Models’ Performance across Different Literal-based ICL Tasks.
@@
14M 31M 70M 160M 410M 1B 1.4B 2.8B 6.9B 12B 3.2-1B 3.2-3B 3.1-8B 3.1-8B-Inst
||||P|||| ||||R|||| Literal Sequence Copying (lsc)
5 5 2.1 1.7 6.5 45.4 96.3 95.8 93.0 95.1 97.7 98.4 99.7 100.0 100.0 99.9
||||P|||| ||||R|||| ||||G|||| Literal Sequence Copying with Random Gaps (lscg)
10 10 5 1.5 0.9 0.6 17.9 64.9 71.9 79.2 86.5 74.1 84.4 96.6 97.8 97.3 93.2
10 10 10 1.1 0.9 0.2 8.7 45.2 55.6 61.9 76.1 58.9 70.7 88.7 93.7 93.7 73.0
20 20 5 1.2 1.0 0.1 8.9 61.2 79.1 78.6 87.6 65.7 80.2 95.6 97.9 96.6 95.8
20 20 10 0.4 0.2 0.1 5.6 43.2 67.9 70.3 73.4 50.4 64.8 92.7 96.3 95.5 85.3
||||S|||| $i$ Word Index (wi)
4 0 0.0 0.1 0.1 21.4 61.9 72.3 84.7 85.5 68.0 77.8 34.9 47.3 87.8 90.4
4 1 0.0 0.1 0.0 2.8 12.8 15.7 21.5 18.4 24.2 22.4 12.4 26.4 40.0 30.0
4 2 0.0 0.0 0.0 2.4 23.3 25.9 27.7 23.6 29.9 29.7 23.6 28.6 33.8 23.3
4 3 0.0 0.0 0.0 16.2 62.6 78.0 78.9 89.3 84.9 90.1 93.5 97.0 99.1 98.6
||||F|||| ||||L|||| ||||D|||| Word Content (wc)
2 2 0 6.5 12.7 25.5 55.2 97.5 96.7 97.5 96.4 99.4 99.4 99.3 99.0 100.0 100.0
2 2 1 4.4 10.3 14.3 46.2 92.4 95.5 94.9 91.8 98.3 98.6 97.9 97.8 100.0 99.7
2 2 2 4.7 10.0 13.7 40.4 87.6 91.1 88.9 88.0 93.3 95.0 95.5 95.9 99.1 99.4
2 2 5 5.8 8.1 16.0 40.9 75.1 77.9 75.7 71.5 81.5 82.3 77.5 75.0 90.7 93.9
2 2 10 5.6 9.9 18.6 37.1 65.7 65.0 66.9 60.5 68.3 66.9 57.4 57.8 71.8 75.3
@@ 4.1 ICL is not Pure Memorization
Table 2 shows an excerpt of the Pythia models’ performance on the pattern-based tasks. When we zoom in to the lsc task, we can observe that every model with a size larger than 400M parameters achieves near-perfect performance. [^3] Full results from more settings and on the LLaMA models are presented in Appendix B, which also supports these findings. We note that all tokens used in the task are randomly sampled, and therefore, this task is unseen to LMs which are pre-trained on natural language text. Thus, if the model only relies on token co-occurrence statistics and does not possess any capability to generalize in the algorithmic and structural manner we detail above, the model cannot achieve the high performance demonstrated. Therefore, consistent with Golchin et al. (2024), we confirm that certain aspects of ICL are independent of memorization. While some instances of ICL may arise from memorization, this pattern-matching and copying capability clearly goes beyond mere memorization. Our findings, in this context of large LMs trained on natural language corpora, reproduce Olsson et al. (2022)’s findings on toy Transformer models. However, we emphasize that this does not imply that ICL is a symbolic algorithm that works irrespective of the task structure or token statistics. To evaluate this, we next explore the effects of unigram token frequency and task configurations—designed to loosely control the notion of the complexity or difficulty of the tasks—on the ICL performance.
@@ 4.2 ICL Performance Strongly Correlates with Token Frequency
Does this ability of LMs to generalize beyond memorization mean the models have learned to carry out some symbolic algorithm, as Olsson et al. (2022) originally posited? Here, we present evidence that LMs’ pattern-repeating behavior is still dependent on token co-occurrence statistics. Specifically, LMs’ pattern-repeating performance strongly correlates with the frequency of the tokens that make up the patterns in the task.
Nevertheless, calculating the frequency of individual tokens in web-scale pre-training data requires a prohibitively large amount of computation resource, and many LMs—such as the GPT and LLaMA families—have not disclosed their pre-training datasets. Therefore, we apply a novel method that uses the tokenizer index as a proxy for the unigram frequency of a token in the tokenized corpus. Most modern LMs, including Pythia and LLaMA (the two types of models used in this work), rely on tokenizers based on the byte pair encoding (BPE; Sennrich et al., 2016) algorithm, where the most common contiguous character sequences are merged into tokens first. As such, tokens with smaller indices (merged earlier in the process) are likely to be more frequent than those with larger indices. [^4] The BPE tokenizer is often trained on a different dataset, and the token indices do not map directly to unigram frequencies, but this remains a reasonable estimate of token frequency. In this experiment, we test whether LMs perform differently on tasks involving tokens with different indices.
@@
Pearson $r$0.8511 -0.6722 -0.6540 -0.9126 -0.7392 -0.6571 -0.7010 -0.7958 -0.6967 -0.7591
$p$-value 9.44e-15 1.23e-07 3.47e-07 6.99e-20 1.31e-09 2.93e-07 2.03e-08 8.26e-12 2.71e-08 2.61e-10
Here, we construct the lsc task [^5] We set both the pattern length and the gap length to 10. Other settings show similar results. using two types of tokens. First, we use the same frequent, full-word tokens sampled from the Brown corpus as in the previous experiment. Second, we divide the tokenizer’s index range into 1000-token intervals (e.g., 10k-11k and 30k-31k) and directly sample tokens from those intervals. Figure 1 shows how the LMs perform on the lsc task with tokens sampled from different ranges. Overall, we observe a general trend: models perform better on more frequent tokens (lower token indices) and worse on
less frequent ones (higher token indices). This trend is statistically significant. In Table 3, we run a Pearson correlation test and find that there is a strong, negative correlation between the LM’s lsc performance and the token indices across across every model size. Moreover, all LMs perform better on the full-word token control group, achieving 98.6–99.9% lsc accuracy, compared to every tokenizer index range. Results for Llama models are shown in Appendix C, and they support the same conclusion.
@@ 5.1 In-distribution ICL Develops First, and the In- and Off-distribution Gap Stabilizes Over Time
We start by investigating the early periods of the pre-training process. We find that LMs’ in- and off-distribution generalization capabilities develop at different paces. Figure 4 shows how Pythia-12b develops the lsc. Specifically, in-distribution generalization ability develops earlier and faster, as we observe that lsc with more frequent tokens improves more quickly compared to those with less frequent tokens. From step 1 to 512, the model shows no meaningful lsc capability; i.e., performance remains close to 0 across all these checkpoints. However, at earlier steps (e.g., steps 64 and 512 in Figure 4(a)), the model achieves some accuracy (0.1% to 0.8%) on the most frequent index ranges, while showing no accuracy on the rest. At step 1000, some meaningful lsc capability begins to appear, with performance following a long-tailed distribution. By steps 2000 to 4000, the disparity in performance across token index ranges diminishes, though it does not vanish entirely. In summary, these observations indicate that in-distribution generalization develops earlier and more rapidly during pre-training compared to the harder off-distribution generalization ability.
Moreover, we evaluate using continuous metrics (particularly, log probability [^8] All of our tasks have exactly one correct answer token, and we follow Schaeffer et al.’s (2023) suggestion to apply the per-token log probability measure for this task setup.) in Figure 5. As Schaeffer et al. (2023) pointed out, accuracy—as a discrete measure—can obscure gradual improvements, fail to
Figure 5: Early development of Pythia-12b’s lsc ability measured in log probability, a continuous metric. The result supports the same findings.
reflect meaningful changes in model behavior, and even lead to misleading conclusions. In our case of lsc, using a continuous metric does reveal the same dispersion of in- and off-domain development. The model’s performance shows a clear stratification by token frequencies (tokenizer indices), with in-distribution lsc ability on more frequent tokens developing earlier and faster, and vice versa for off-distribution ability on rare ones.
@@ 5.2 ICL Competence for Different Task Configurations Develops in Lock-step
Figure 9: Token translation (tt) task performance improves steadily with model size, measured by the best checkpoint’s performance. The trend is also predictable, as it can be modeled using a scaling law of performance with respect to model size ($N$).
@@ 5.3 A Scaling Law: the Development of Complex Aspects of ICL is Gradual and Predictable
Whether LM abilities develop abruptly and unpredictably, as famously proposed by Wei et al. (2022), or whether the process follows a more predictable pattern (Lu et al., 2024; Schaeffer et al., 2023), remains a topic of heated debate in the interpretability and language modelling community. Our investigation of more complex tasks suggests that ICL competence—an essential component of LM abilities—develops gradually and predictably, following a describable scaling law. The rapid development of ICL competence observed by Olsson et al. (2022), which led to an abrupt phase change, however, is only seen in the easier synthetic tasks in our study. Further insights could be gained by extending our analysis to more fine-grained checkpointing intervals, particularly between steps 512 and 1000. However, to the best of our knowledge, no such publicly available repository exists beyond Pythia. Thus, we leave this for future work.

### v3

@@ 3 Experimental Setup: Tasks, Data and Models
Previous work primarily defines ICL as the ability to learn an input–-label mapping based on explicit in-context samples, with the only exception being Elhage et al. (2021), who investigated literal sequence copying as a precursor to ICL. In this work, we argue that any next-token prediction that requires the model to recover a token distribution from the context and utilize it (with or without another token distribution memorized from the pre-training) qualifies as ICL—a notion of ICL very similar to Xie et al. (2021). This definition allows us to define pattern-matching tasks similar to Olsson et al. (2022), standard ICL examples of input-label pairs, as well as counterfactual reasoning with in-context facts under the broad umbrella of ICL.
@@
To the best of our knowledge, all prior works evaluating ICL (Conneau & Kiela, 2018; Hernandez et al., 2023; Todd et al., 2023) have used tasks that require some form of factual, common-sense, or linguistic knowledge and therefore may be subject to data contamination. Since these web-scale training datasets often comprise billions or even trillions of tokens, we cannot completely rule out the possibility—even when the pre-training dataset is fully accessible (e.g., Biderman et al., 2023; Groeneveld et al., 2024). Additionally, with real-world, natural-language tasks, it is practically impossible to control task configurations (e.g., feature sparsity, length dependence, token frequency statistics, reliance on parametrized knowledge, etc.), and as a result, one cannot reliably correlate model scale (training time and parameter size) with specific task characteristics.
Proposed task suit. To address the limitations of using existing benchmark data as explained earlier, in Table 1, we propose a novel suite of tasks constructed from various patterns using randomly sampled tokens, including literal sequence copying (lsc), literal sequence copying with random gaps (lscg), word content (wc) and word index (wi). These tasks satisfy the notion of ICL as recovering and inferring from token distributions in-context. Since these tasks appear as sequences of random tokens, it is functionally impossible for them to exist in the pre-training dataset. Therefore, if an LM can perform these tasks, it cannot be due to memorization—some form of pattern generalization must be at play. Moreover, we introduce two tasks focused on associating and overriding parameterized information using the ICL pattern identification capability: we use token translation [^2] For the tt task, we sample 200 common words across four lanugages: English, German, Spanish, and Italian. The sampling process is detailed in Appendix A. (tt) to investigate how patterns are combined with parameterized information, and counterfactuals (cf) to observe how this parameterized information can be overridden by counterfactual prompt settings. Since these tasks require specific parameterized information, we cannot guarantee that it does not appear in the pre-training dataset; however, this is acceptable, as we already include pattern-based tasks that do not rely on such information and serve as a cleaner test of ICL capabilities. We assume that they are “harder” than the pattern-based tasks, as they involve an additional layer of combining or overriding information.
@@
Table 2: Selected Models’ Performance across our proposed tasks.
@@
14M 31M 70M 160M 410M 1B 1.4B 2.8B 6.9B 12B 3.2-1B 3.2-1B-I 3.2-3B 3.2-3B-I 3.1-8B 3.1-8B-I
||P|| ||R|| Literal Sequence Copying (lsc)
5 5 2.1 1.7 6.5 45.4 96.3 95.8 93.0 95.1 97.7 98.4 99.7 98.4 100.0 99.3 100.0 99.9
10 10 3.0 1.6 4.8 51.8 99.3 98.6 98.7 99.7 99.8 99.9 99.6 100.0 100.0 99.8 100.0 100.0
20 20 2.1 7.2 2.4 60.0 99.2 99.7 99.7 99.8 99.8 100.0 99.8 99.9 100.0 99.9 100.0 100.0
||P|| ||R|| ||G|| Literal Sequence Copying with Random Gaps (lscg)
10 10 0 2.4 2.6 4.0 55.4 98.4 99.3 99.0 99.6 99.8 99.9 100.0 99.9 100.0 99.9 100.0 100.0
10 10 5 1.5 0.9 0.6 17.9 64.9 71.9 79.2 86.5 74.1 84.4 96.6 92.0 97.8 87.0 97.3 93.2
10 10 10 1.1 0.9 0.2 8.7 45.2 55.6 61.9 76.1 58.9 70.7 88.7 71.0 93.7 55.4 93.7 73.0
20 20 0 1.8 5.9 1.9 55.7 99.6 99.6 99.7 99.9 99.9 100.0 100.0 100.0 100.0 99.8 100.0 100.0
20 20 5 1.2 1.0 0.1 8.9 61.2 79.1 78.6 87.6 65.7 80.2 95.6 93.5 97.9 91.7 96.6 95.8
20 20 10 0.4 0.2 0.1 5.6 43.2 67.9 70.3 73.4 50.4 64.8 92.7 87.0 96.3 75.2 95.5 85.3
||S|| $i$ Word Index (wi)
4 0 0.0 0.1 0.1 21.4 61.9 72.3 84.7 85.5 68.0 77.8 34.9 88.6 47.3 70.0 87.8 90.4
4 1 0.0 0.1 0.0 2.8 12.8 15.7 21.5 18.4 24.2 22.4 12.4 19.7 26.4 19.5 40.0 30.0
4 2 0.0 0.0 0.0 2.4 23.3 25.9 27.7 23.6 29.9 29.7 23.6 18.8 28.6 15.4 33.8 23.3
4 3 0.0 0.0 0.0 16.2 62.6 78.0 78.9 89.3 84.9 90.1 93.5 96.8 97.0 98.3 99.1 98.6
||F|| ||L|| ||D|| Word Content (wc)
2 2 0 6.5 12.7 25.5 55.2 97.5 96.7 97.5 96.4 99.4 99.4 99.3 93.9 99.0 97.9 100.0 100.0
2 2 1 4.4 10.3 14.3 46.2 92.4 95.5 94.9 91.8 98.3 98.6 97.9 88.3 90.7 97.8 100.0 99.7
2 2 2 4.7 10.0 13.7 40.4 87.6 91.1 88.9 88.0 93.3 95.0 95.5 82.8 84.2 95.9 99.1 99.4
2 2 5 5.8 8.1 16.0 40.9 75.1 77.9 75.7 71.5 81.5 82.3 77.5 73.0 69.9 75.0 90.7 93.9
2 2 10 5.6 9.9 18.6 37.1 65.7 65.0 66.9 60.5 68.3 66.9 57.4 59.5 58.4 57.8 71.8 75.3
Setting Token Translation (tt)
en $\rightarrow$ de 0.1 0.1 0.7 2.9 22.2 32.0 54.9 70.9 77.6 82.5 76.40 79.00 87.90 85.50 90.20 89.70
en $\rightarrow$ fr 0.1 0.1 0.8 5.5 29.3 36.2 63.2 75.2 85.5 85.1 83.30 79.00 91.30 89.20 93.10 93.20
en $\rightarrow$ es 0.5 0.2 0.4 4.3 30.8 40.7 66.7 78.3 82.2 86.9 84.80 77.50 87.90 85.20 91.80 91.00
en $\rightarrow$ it 0.2 0.0 0.7 4.0 20.3 33.3 57.3 74.7 80.0 84.9 82.90 74.30 89.00 85.90 92.90 93.90
Counterfactual (cf)
cf 0.0 0.0 0.1 2.5 7.1 33.3 22.9 72.6 85.0 87.9 73.0 83.0 95.9 95.9 98.4 97.1
@@ 4.1 ICL is not Pure Memorization
Table 2 shows an excerpt of the Pythia models’ performance on the pattern-based tasks. When we zoom in to the lsc task, we can observe that every model with a size larger than 400M parameters achieves near-perfect performance. [^3] Full results from more settings and on the LLaMA models are presented in Appendix B, which also supports these findings. We note that all tokens used in the task are randomly sampled, and therefore, this task is unseen to LMs which are pre-trained on natural language text. Thus, if the model only relies on token co-occurrence statistics and does not possess any capability to generalize in the algorithmic and structural manner we detail above, the model cannot achieve the high performance demonstrated. Therefore, consistant with Golchin et al. (2024), we confirm that certain aspects of ICL are independent of memorization. While some instances of ICL may arise from memorization, this pattern-matching and copying capability clearly goes beyond mere memorization. Our findings, in this context of large LMs trained on natural language corpora, reproduce Olsson et al. (2022)’s findings on toy Transformer models. However, we emphasize that this does not imply that ICL is a symbolic algorithm that works irrespective of the task structure or token statistics. To evaluate this, we next explore the effects of unigram token frequency and task configurations—designed to loosely control the notion of the complexity or difficulty of the tasks—on the ICL performance.
@@ 4.2 ICL Performance Strongly Correlates with Token Frequency
But, does this ability of LMs to generalize beyond memorization mean the models have learned to carry out some symbolic algorithm, as Olsson et al. (2022) originally posited? Here, we present evidence that LMs’ pattern-repeating behavior is still dependent on token co-occurrence statistics. Specifically, LMs’ pattern-repeating performance strongly correlates with the frequency of the tokens that make up the patterns in the task.
Nevertheless, calculating the frequency of individual tokens in web-scale pre-training data requires a prohibitively large amount of computation resource, and many LMs—such as the GPT and LLaMA families—have not disclosed their pre-training datasets. Therefore, we apply a novel method that uses the tokenizer index as a proxy for the unigram frequency of a token in the tokenized corpus. Most modern LMs, including Pythia and LLaMA (the two types of models used in this work), rely on tokenizers based on the byte pair encoding (BPE; Sennrich et al., 2016) algorithm, where the most common contiguous character sequences are merged into tokens first. As such, tokens with smaller indices (merged earlier in the process) are likely to be more frequent than those with larger indices. [^4] The BPE tokenizer is often trained on a different dataset, and the token indices do not map directly to unigram frequencies. Moreover, “open-weight” LMs such as Llama do not release their pre-training datasets, making direct comparison practically impossible. Therefore, we sampled 5,000 tokens from the Pythia tokenizer, counted their token frequencies on 250,000 documents from the Pile, and computed the correlation between tokenizer index and token count. We observed a Spearman r=0.746r=-0.746 and a Pearson r=0.701r=-0.701, both with $p<10^{-100}$. This suggests that the tokenizer index is a reasonable and effective approximation of unigram token frequency statistics. In this experiment, we test whether LMs perform differently
on tasks involving tokens with different indices.
@@
Pearson $r$ -0.8511 -0.6722 -0.6540 -0.9126 -0.7392 -0.6571 -0.7010 -0.7958 -0.6967 -0.7591
pp-value 9.44e-15 1.23e-07 3.47e-07 6.99e-20 1.31e-09 2.93e-07 2.03e-08 8.26e-12 2.71e-08 2.61e-10
Here, we construct the lsc task [^5] We set both the pattern length and the gap length to 10. Other settings show similar results. using two types of tokens. First, we use the same frequent, full-word tokens sampled from the Brown corpus as in the previous experiment. Second, we divide the tokenizer’s index range into 1000-token intervals (e.g., 10k–11k and 30k–31k) and directly sample tokens from those intervals. Figure 1 shows how the LMs perform on the lsc task with tokens sampled from different ranges. Overall, we observe a general trend: models perform better on more frequent tokens (lower token indices) and worse on less frequent ones (higher token indices). This trend is statistically significant. In Table 3, we run a Pearson correlation test and find that there is a strong, negative correlation between the LM’s lsc performance and the token indices across across every model size. Moreover, all LMs perform better on the full-word token control group, achieving 98.6–99.9% lsc accuracy, compared to every tokenizer index range. Results for Llama models are shown in Appendix C, results for other literal-based ICL tasks (lscg, wc, wi) are shown in Appendix D and they support the same conclusion.
@@ 4.3 ICL Performance Varies with Task Configuration Choices
In sum, we argue that a true “symbolic algorithm” should be robust to minor changes in task structure. However, we find that task performance is related to three simple factors: the gap length in the lscg task, the sequence length in the wi task, and the number of distractors or features in the wc task.
@@ 5.1 In-distribution ICL Develops First, and the In- and Off-distribution Gap Stabilizes Over Time
We start by investigating the early periods of the pre-training process. We find that LMs’ in- and off-distribution generalization capabilities develop at different paces. Figure 4 shows how Pythia-12b develops the lsc. Specifically, in-distribution generalization ability develops earlier and faster, as we observe that lsc with more frequent tokens improves more quickly compared to those with less frequent tokens. From step 1 to 512, the model shows no meaningful lsc capability; i.e., performance remains close to 0 across all these checkpoints. However, at earlier steps (e.g., steps 64 and 512 in Figure 4(a)), the model achieves some accuracy (0.1% to 0.8%) on the most frequent index ranges, while showing no accuracy on the rest. At step 1000, some meaningful lsc capability begins to appear, with performance following a long-tailed distribution. By steps 2000 to 4000, the disparity in performance across token index ranges diminishes, though it does not vanish entirely. In summary, these observations indicate that in-distribution generalization develops earlier and more rapidly during pre-training compared to the harder off-distribution
Figure 5: The early development of Pythia-12b’s lsc ability measured in log probability, a continuous metric. The result supports the same findings.
generalization ability.
Moreover, we evaluate using continuous metrics (particularly, log probability [^8] All of our tasks have exactly one correct answer token, and we follow Schaeffer et al.’s (2023) suggestion to apply the per-token log probability measure for this task setup.) in Figure 5. As Schaeffer et al. (2023) pointed out, accuracy—as a discrete measure—can obscure gradual improvements, fail to reflect meaningful changes in model behavior, and even lead to misleading conclusions. In our case of lsc, using a continuous metric does reveal the same dispersion of in- and off-domain development. The model’s performance shows a clear stratification by token frequencies (tokenizer indices), with in-distribution lsc ability on more frequent tokens developing earlier and faster, and vice versa for off-distribution ability on rare ones.
@@ 5.2 ICL Competence for Different Task Configurations Develops in Lock-step
Figure 9: Token translation ($t$) task performance improves steadily with model size, measured by the best checkpoint’s performance. The trend is also predictable, as it can be modeled using a scaling law of performance with respect to model size (NN).
@@ 5.3 A Scaling Law: the Development of Complex Aspects of ICL is Gradual and Predictable
Whether LM abilities develop abruptly and unpredictably, as famously proposed by Wei et al. (2022), or whether the process follows a more predictable pattern (Lu et al., 2024; Schaeffer et al., 2023), remains a topic of heated debate in the interpretability and language modeling community. Our investigation of more complex tasks suggests that ICL competence—an essential component of LM abilities—develops gradually and predictably, following a describable scaling law. The rapid development of ICL competence observed by Olsson et al. (2022), which led to an abrupt phase change, however, is only seen in the easier synthetic tasks in our study. Further insights could be gained by extending our analysis to more fine-grained checkpointing intervals, particularly between steps 512 and 1000. However, to the best of our knowledge, no such publicly available repository exists beyond Pythia. Thus, we leave this for future work.